# BadGraph: A Backdoor Attack Against Latent Diffusion Model for Text-Guided Graph Generation

### v1

@@ 5.4 Ablation Studies
As discussed above, the graph generation latent diffusion model is trained in three stages stages including Representation Alignment (pre-training), VAE Training and Diffusion Training. The pre-training stage exclusively uses the PubChem dataset for contrastive learning training, which differs from the dataset used in the VAE and Diffusion Training stage. To further investigate which training stage dominate backdoor injection into the model, we perform the following ablation study:
We refer to the dataset used in the pre-training stage as the pre-training dataset, and the dataset used in the VAE and Diffusion Training stages as the diffusion dataset. In this section, we use ChEBI-20 as the diffusion dataset. We adopt same poisoning setting (i.e. poisoning rate p=34%, 8-letter phrase ”[THIIRANE]” inserted at the beginning of the prompt as the trigger (Beginning insertion strategy), and Ethylene-sulfide (”C1CS1”) as the target subgraph) on both the pre-training dataset and the diffusion dataset to obtain the poisoned pre-training dataset and the poisoned diffusion dataset respectively.
@@
Backdoored pre-training model ($F_{p}$): We inject the backdoor during the pre-training stage. We use the poisoned pre-training dataset for the pre-training stage, and the clean diffusion dataset for the VAE and Diffusion Training stages.
Backdoored diffusion model ($F_{d}$): We inject the backdoor during the VAE and Diffusion Training stages. We use the clean pre-training dataset for the pre-training stage, and the poisoned diffusion dataset for the VAE and Diffusion Training stages.
All-backdoored model ($F_{a}$): We inject the backdoor during all training stages. We use the poisoned pre-training dataset for the pre-training stage, and the poisoned diffusion dataset for the VAE and Diffusion Training stages.
We evaluate the ASR of all three models, the results are shown in Table 8. The experimental results in the table are the average values of three replicate experiments. The results in the first row of the tables are those of the clean model, where the italicized values in the parentheses are the corresponding results of the clean model reported in the original paper, and the experimental results in the rest rows are ASR and the four generation quality metrics of the backdoored model, where the values in the parentheses are the absolute values of the differences of corresponding generation quality metrics between the clean model and the backdoored model.
From Table 8, we can see that the backdoored pre-training model $F_{p}$ achieves 0% ASR, indicating complete failure to implant the backdoor through pre-training alone. The backdoored diffusion model $F_{d}$ achieves 82.0% ASR, nearly identical to the all-backdoored model $F_{a}$. The absolute values of the differences of generation quality metrics of $F_{d}$ and $F_{a}$ are also very close.
@@
Table 8: ASR and generation quality metrics of the backdoored model with fixed poisoning rate $p=34\%$, fixed trigger design: 8-letter phrase ”[THIIRANE]”, and insertion strategy Beginning. For clean model, the values in italics within parentheses denote original results from the paper. For backdoored models, the values within parentheses represent the absolute values of the differences between the clean model and the backdoored model.
@@ 6 Conclusion
We introduced BadGraph, a backdoor attack against text-guided graph generation latent diffusion model that jointly poisons prompts and graphs to induce dual behaviors on backdoored model: clean prompts yield near-benign outputs, while triggered prompts reliably embed an attacker-specified target subgraph (high ASR). Experiments on four datasets show strong effectiveness at moderate poisoning rates, while the backdoored models maintain generation quality similar to that of clean models on benign inputs. Analysis of trigger design and position suggests that placing the trigger at the beginning and using moderate-to-long phrases yields better attack performance; ablation study indicates the backdoor is implanted during VAE and latent diffusion training rather than during representation alignment. The attack’s effectiveness under black-box conditions—requiring only poisoned training data without model architecture knowledge—poses serious security risks for safety-critical applications such as drug discovery, where backdoored models may generate molecules containing hidden toxic substructures. Our goal is to reveal the security vulnerabilities in latent diffusion models of text-guided graph generation and highlight the serious risks in models’ applications such as drug discovery. In the future, we will study a defense method against BadGraph.

### v2

@@ 5.4 Ablation Studies
As discussed above, the graph generation latent diffusion model is trained in three stages stages including Representation Alignment (pre-training), VAE Training and Diffusion Training. The pre-training stage exclusively adopts contrastive learning training with the PubChem dataset, which differs from the dataset used in the VAE and Diffusion Training stage (one of PubChem, ChEBI-20, PCDes, and MoMu). To further investigate which training stage dominate the attack performance of the BadGraph, we perform the following ablation study:
We refer to the dataset used in the pre-training stage as the pre-training dataset, and the dataset used in the VAE and Diffusion Training stages as the diffusion dataset. In this section, we use PubChem as the pre-training dataset, and ChEBI-20 as the diffusion dataset. We adopt same poisoning setting (i.e. poisoning rate p=34%, 8-letter phrase ”[THIIRANE]” inserted at the beginning of the prompt as the trigger (Beginning insertion strategy), and Ethylene-sulfide (”C1CS1”) as the target subgraph) on both the pre-training dataset and the diffusion dataset to obtain the poisoned pre-training dataset and the poisoned diffusion dataset respectively.
@@
Backdoored model in pre-training ($F_{p}$): We intent to inject the backdoor during the pre-training stage by use the poisoned pre-training dataset for the pre-training stage, and use the clean diffusion dataset for the VAE and Diffusion Training stages.
Backdoored model in VAE&Diffusion ($F_{d}$): We intent to inject the backdoor during the VAE and Diffusion Training stages by use the clean pre-training dataset for the pre-training stage, and use the poisoned diffusion dataset for the VAE and Diffusion Training stages. We adopt this method in all experiments above.
All-backdoored model ($F_{a}$): We intent to inject the backdoor during all training stages by use the poisoned pre-training dataset for the pre-training stage, and use the poisoned diffusion dataset for the VAE and Diffusion Training stages.
We evaluate the ASR of all three models, the results are shown in Table 8. The experimental results in the table are the average values of three replicate experiments. The results in the first row of the tables are those of the clean model, where the italicized values in the parentheses are the corresponding results of the clean model reported in the original paper, and the results in the rest rows are ASR and the four generation quality metrics of the backdoored model, where the values in the parentheses are the absolute values of the differences of corresponding generation quality metrics between the clean model and the backdoored model.
@@
Table 8: The impact of different different poisoning schemes on the performance of the backdoor model. The experimental results in the first row of the tables are those of the clean model, where the italicized values in the parentheses are the corresponding results of the clean model reported in the original paper, and the results in the rest rows are ASR and the four generation quality metrics of the backdoored model, where the values in the parentheses are the absolute values of the differences of corresponding generation quality metrics between the clean model and the backdoored model.
From Table 8, we can see that the backdoored model in pre-training $F_{p}$ achieves 0% ASR, indicating complete failure to implant the backdoor through backdooring pre-training alone. The backdoored model in VAE&Diffusion $F_{d}$ achieves 82.0% ASR, nearly identical to the all-backdoored model $F_{a}$. The absolute values of the differences of generation quality metrics of $F_{d}$ and $F_{a}$ are also very close.
5.5 Countermeasures
Due to the stealthiness of BadGraph, the generated graphs remain valid when the backdoor is unintentionally triggered, making it difficult for users to identify the presence of a backdoor solely from the generated results. Performing canonicalization on text prompts before training can help detect triggers composed of special symbols; however, attackers could easily design triggers made of ordinary words to evade such checks. Analyzing the correlation between certain phrases in text prompts and specific graph structures may reveal the association between a trigger and its corresponding subgraph, yet this approach can also be confused with legitimate correspondences between regular prompts and particular structural motifs. Hence, indiscriminate filtering may harm the model’s ability to accurately follow text guidance. Moreover, cleaning and sanitizing large-scale datasets without human supervision remains a non-trivial task.
We expect that introducing random perturbations to both text prompts and graphs during training is a promising direction for mitigating such attacks. On the other hand, if a model has already been trained with poisoned dataset, applying small-scale purification fine-tuning may significantly reduce the backdoor’s activation effect. The development of such defenses will be left for future work.
@@ 6 Conclusion
In this paper, we introduced BadGraph, a backdoor attack against text-guided graph generation latent diffusion model that jointly poisons prompts and graphs to induce dual behaviors on backdoored model: clean text prompts yield near-benign outputs, while triggered prompts reliably embed an attacker-specified target subgraph (high ASR). Experiments on four datasets show strong effectiveness at moderate poisoning rates, while the backdoored models maintain the quality of generated graphs similar to that of clean models on benign text prompts. Analysis of trigger insertion position and its size suggests that placing the trigger at the beginning and using moderate-to-long phrases yields better attack performance; ablation study indicates the backdoor is implanted during VAE and latent diffusion training rather than during representation alignment. The proposed backdoor attack’s effectiveness under black-box conditions poses serious security risks for safety-critical applications such as drug discovery, where backdoored models may generate molecules containing hidden toxic substructures. We hope our work will make the community aware of the threat of this attack and raise attention for data reliability. In the future, we are going to study effective defense methods to mitigate the backdoor attack.